Find 3 consecutive integers such that the sum of twice the smallest and 3 times the largest is 126.
Answer by checkley77(12569)

LET X, X+1 & X+2 BE THE THREE INTEGERS.
2X+3(X+2)=126
2X+3X+6=126
5X=126-6
5X=120
X=120/5
X=24 ANSWER FOR THE SMALLEST INTEGER.
24+1=25 IS THE MIDDLE INTEGER.
24+2=26 FOR THE LARGEST INTEGER.
PROOF:
2*24+3*26=126
48+78=126
126=126